Write in simplest form:
9-b^2
---------
b^2-b-6
I have four answers to choose form but do not come up with any of them each time, here they are:
A) -b+3
------
b+2
B) -b+3
------
b-2
C) b+3
------
b+2
D) -b-3
------
b+2
Thank you for the help!!!

Answer by ptaylor(2198)

Write in simplest form:
9-b^2
---------
b^2-b-6

NUMERATOR: 9-b^2=(3+b)(3-b)=-(b-3)(b+3)
DENOMINATOR: b^2-b-6=(b-3)(b+2)
Now we put 'em back together
9-b^2
--------- =
b^2-b-6



-(b-3)(b+3)
----------- =
(b-3)(b+2)




-(b+3)
------=
(b+2)





-b-3
----
b+2

LOOKS LIKE D IS THE CORRECT ANSWER

Hope this helps-----ptaylor